OpenStudy (anonymous):

7^.34 how do you get the answer when you power a decimal?

OpenStudy (anonymous):

we could change the decimal to a fraction, do you know how fractional powers?

OpenStudy (anonymous):

how fractional powers work*

OpenStudy (anonymous):

34/100

OpenStudy (anonymous):

yep, and simplified: 17/50

OpenStudy (anonymous):

then whats the next step?

OpenStudy (anonymous):

so: \[7^{0.34} = 7^{\frac{17}{50}} = (7^{17})^{\frac{1}{50}}=\sqrt[50]{7^{17}}\]

OpenStudy (anonymous):

well you see i knew how to do that ... but i don't know how to apply this to my problem which is f(x) = 15,000x^.34 f(7)

OpenStudy (anonymous):

can you use a calculator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what happens if you just try to type in 15000*(7^(0.34)) ?

OpenStudy (anonymous):

well its easy on the calculator but i want to learn how to do it not just depend on a calculator... do you know how to solve it? manually?

OpenStudy (anonymous):

yes, like above first we know 7^(0.34) = 7^(17/50) so we calculate 7^15 , which comes to a very big number about 4.74756151 × 10^12 then we take the 50th root of that number which would take us ages, and is really why we use a calculator alternatively if you have log tables and know how to use them then that would help you get a quick estimate

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

happy to help :)